# Evaluating LLMs on Sequential API Call Through Automated Test Generation

### v1

@@ Abstract
By integrating tools from external APIs, Large Language Models (LLMs) have expanded their promising capabilities in a diverse spectrum of complex real-world tasks. However, testing, evaluation, and analysis of LLM tool use remain in their early stages. Most existing benchmarks rely on manually collected test cases, many of which cannot be automatically checked for semantic correctness and instead depend on static methods such as string matching. Additionally, these benchmarks often overlook the complex interactions that occur between sequential API calls, which are common in real-world applications. To fill the gap, in this paper, we introduce StateGen, an automated framework designed to generate diverse coding tasks involving sequential API interactions. StateGen combines state-machine-based API constraint solving and validation, energy-based sampling, and control-flow injection to generate executable programs. These programs are then translated into human-like natural language task descriptions through a collaboration of two LLM agents. Utilizing StateGen, we construct StateEval, a benchmark encompassing 120 verified test cases spanning across three representative scenarios: Session Service, Tensor Operation, and ElevenLabs MCP. Experimental results confirm that StateGen can effectively generate challenging and realistic API-oriented tasks, highlighting areas for improvement in current LLMs incorporating APIs.
Index Terms:
Large Language Models, Code Generation, Tool Use, Sequential Function Call, Benchmark, Fuzzing
@@ I Introduction
The meteoric advancements in Large Language Models (LLMs) have enabled significant performance leaps across domains [1, 2, 3] and have transformed many areas of software engineering [4, 5, 6, 7, 8, 9]. Building on this progress, stakeholders have found that integrating tool usage with LLM-driven systems can further extend their capabilities [10, 11, 12, 13], enabling tasks such as numerical computation [14] and access to up-to-date information [15] that are otherwise difficult for LLMs to perform on their own.
Despite the promising progress of this tool integration paradigm, using LLMs to perform designated tasks with external APIs remains a practical yet challenging problem. In real-world software development, practitioners typically start by analyzing and decomposing the task requirements and then understanding the functionalities and usage of available APIs from their documentation [16]. Eventually, multiple APIs need to be managed in an appropriate sequence, with correctly determined inputs and parameters, to achieve the intended functionalities [17, 18]. Such a multi-step development pipeline demands advanced abilities in reasoning, management, planning, and tool calling, posing significant challenges for LLMs, which act as the central control unit to coordinate these processes. As illustrated in Figure 1, converting a given text into speech with a specific style demands the LLM to extract key user requirements (i.e., “convert to speech”, “professional feminine voice”) from the input prompt and inspect the available APIs to determine the appropriate ones for execution. Given the numerous sequences of API calls, the LLM must identify the correct one capable of accomplishing the task and manage the parameters and variables accordingly. Considering the complexity involved, it is crucial to understand to what extent LLMs can comprehend the functionalities of various APIs and facilitate sequential API calls to fulfill designated objectives. Therefore, systematic testing and evaluation of such intricate yet realistic scenarios are one of the foundational components for moving the field forward.
@@
Several recent benchmarks [5, 19, 20, 21] have been introduced to systematically evaluate the coding capabilities of LLMs across a variety of contexts. However, most of these efforts primarily focus on general-purpose, small-scale coding tasks, often overlooking scenarios that require reasoning over interdependent API interactions. While some studies [22, 23, 24] have begun to incorporate API-centric tasks from specific libraries, their evaluation scope is typically confined to a narrow set of libraries, thereby limiting the generalizability of their findings to broader or novel contexts.
@@
On the other hand, existing benchmarks tailored for function calling tasks tend to rely on manually curated datasets, which are labor-intensive to construct and difficult to scale to new domains or task types. Moreover, many related works concentrate on addressing daily user requests that involve only a handful of APIs per example rather than tackling more complex programming scenarios that demand sophisticated orchestration of multiple APIs. This gap highlights the need for more systematic and scalable benchmarks that can better reflect the challenges encountered in real-world, API-rich programming environments.
@@
Consequently, a key research question thereupon arises: How can we design an automated test generation approach that systematically evaluates LLMs’ ability in understanding sequential API calls and managing the associated program states?
To address this gap, we propose StateGen and StateEval in this study. By leveraging existing API documentation, StateGen generates diverse, executable programs involving sequential API calls through techniques such as API compatibility checking, energy-based sampling, and control-flow injection. Subsequently, two LLM agents collaborate to translate these programs into coherent natural language instructions, serving as test inputs for the target LLMs. Building upon StateGen, we introduce a new benchmark, StateEval, comprising 120 API-oriented test cases across three representative application scenarios. In summary, this paper makes the following contributions:
@@
  • Manually Verified Benchmark. Using this framework, we implement three scenarios: RESTful API calling (Session Service [25]), tensor manipulation (Tensor Operation [26]), and text-to-speech MCP-based LLM tool calling (ElevenLabs MCP [27]). We generate and manually verify 120 test cases, forming a new benchmark, StateEval, to support future evaluation and advancement of LLM-driven API usage.
@@
II Related Work
In this section, we review related work from two perspectives: (1) benchmarks for evaluating LLMs on API calling tasks, and (2) methods for RESTful API fuzzing that involve sequential trace generation.
II-A Evaluating LLMs with API Call
Substantial efforts have been made by practitioners to assess the capability of LLMs incorporating APIs under various scenarios [12, 28, 29, 30] with natural language instructions. Nevertheless, most existing approaches focus on building benchmarks or performing empirical studies by either manually collecting test cases or extracting them from online forums [13, 10, 31, 32, 33, 34, 35].
Early studies have focused on evaluating LLMs’ ability to select the appropriate tool and invoke it at the right time, leveraging metrics such as accuracy for evaluation [13, 12, 31]. For example, Qin et al. developed a benchmark for tool-augmented LLMs that assesses the model’s capability to choose the correct API and populate its parameters based on user queries [12]. Although these studies often cover a wide range of domains collected in real-life, their tasks typically involve fulfilling specific requests rather than generating complete, executable software programs.
BFCL [10] is another example that provides comprehensive evaluations of LLM performance on tasks involving API calls across various programming languages. A key contribution of BFCL is the introduction of parallel and multiple function calling scenarios within executable environments. However, these scenarios still only involve a limited set of APIs, and the API calls typically lack interdependencies between their parameters.
II-B Automated Testing of RESTful APIs
Automated generation of effective and diverse API call sequences remains a commonly studied yet challenging task. One representative research in this field is RESTful API testing [25, 36, 37, 38, 39, 40, 41, 42, 43, 44], which generally follows a black-box approach and relies on standardized specifications such as OpenAPI [45]. This line of work typically adopts a black-box approach, as the source code is often unavailable in API testing. In such cases, Golmohammadi et al. identified the generation of meaningful API sequences as a primary challenge, which is complicated by inter-API dependencies and underspecified schemas [43].
To address this challenge, techniques for API sequence generation largely fall into two categories. Top-down and bottom-up approaches. For the former, RESTTESTGEN [37] is a representative case that first constructs an overall API call graph (e.g., an Operation Dependency Graph), and then instantiates API calls based on this graph. While providing a holistic view, the efficacy of such methods can be compromised by the quality of the initial graph, which may be derived from flawed or incomplete specifications.
Conversely, bottom-up strategies [36, 46] build sequences incrementally. For example, RESTler [36] extends sequences one call at a time using heuristics, while MOREST [46] utilizes a dynamically updating RESTful-service Property Graph to capture detailed API behaviors and dependencies.
@@ III Approach
III-A Problem Definition
To evaluate LLMs in terms of their ability to generate code involving sequential API usage, we formalize the task as a state-dependent code generation problem.
@@ III-A Problem Definition
Let $S=(s_{1},s_{2},\ldots,s_{n})$ denote the set of all program states, where each $s_{j}$ corresponds to its respective $j$-th state. A given state may represent either observable program status (e.g., locally defined variables) or external elements (e.g., the data items stored in remote databases). Given the natural language instructions specifying the desired tasks for a sequence of API calls, along with documentation that defines the usages of corresponding APIs, the objective for LLMs is to generate code that is syntactically and semantically correct in performing the intended functionalities. The generated code should encompass a sequence of $n$ API calls $O=(o_{1},o_{2},\ldots,o_{n})$, where $o_{j}$ denotes the API call at the $j$-th timestamp and can have side effects on states in $S$.
@@
To evaluate the capabilities of LLMs in addressing such intricate API-oriented coding tasks, we propose StateGen, an automated test case generation framework for LLMs with sequential API calls. The overall workflow is shown in Fig 2. At a high level, StateGen follows a reverse-generation strategy, beginning with generating valid, executable sequences of API calls of a given length (Section III-B). These traces are then used as a foundation to construct more intricate executable programs by incorporating control flows, thereby enhancing diversity and complexity. (Section III-C). Next, StateGen employs an LLM-driven multi-agent translation process (Section III-D) to transform these programs into human-like natural language instructions. To obtain test oracles for automated evaluation, StateGen executes the generated programs in a local environment and records the state transitions throughout program execution (Section III-E). These captured transitions serve as ground truth for subsequent evaluation. The following sections detail the design of each aforementioned step accordingly.
III-B Trace Generation
@@
Figure 2: Workflow overview of StateGen. Trace Generation (Sec III-B) forms the backbone of StateGen, producing valid API sequences through state machines while ensuring diversity with energy-based sampling. Program Generation (Sec III-C) then assembles these traces with appropriate initialization and control flow structures. Finally, Instruction Translation (Sec III-D) employs a multi-agent system to convert the generated programs back into natural language descriptions for evaluation.
@@
Generating valid sequential API calls poses significant technical challenges in this work, as it requires precise modelling of program states and meticulous selection of valid API calls at each step. Traditional approaches that address this problem have been extensively studied in the context of web service testing, e.g., RESTful API testing [25, 36, 37, 38, 40, 46, 43]. However, these methods are difficult to apply directly, as they focus primarily on CRUD operations, which represent only a subset of our applications. Adapting them for general LLM tool calling would require significant additional effort.
@@
Despite this, the core concept remains unchanged. Prior work in RESTful API testing can generally be categorized into top-down [37] and bottom-up [36, 46] strategies. In top-down approaches, the overall API call graph is built in advance, and each API call is instantiated according to the graph. Conversely, bottom-up approaches construct the API sequence progressively by adding one call at a time. In this work, we adopt a bottom-up strategy, allowing us to employ a “lazy evaluation” strategy for determining API inputs and parameters. Specifically, we compute only the valid and relevant search space at each step of API selection rather than pre-solving constraints for the entire API call graph, which otherwise would demand substantial computational efforts on optimization to attain practical efficiency [47].
@@
$[o:\mathbb{S}\times\mathbb{P}\rightarrow\mathbb{S}],\quad\forall o\in\mathbb{O}$ (1)
@@
1:$n\in\mathbb{N}^{+}$triangleright Target API call count
2:$O$triangleright Generated trace of API calls
@@
4:     $O\leftarrow\emptyset$, $S\leftarrow RandomInit$triangleright Initialization
5:     $\hat{o}\leftarrow\bot$triangleright Initialize previous transition to None
@@
7:         $\mathcal{A}\leftarrow\{o\in\mathcal{O}\mid\text{valid}(o,S,O)\}$triangleright Obtain transitions
8:         $\mathcal{E}\leftarrow\left\{\frac{1}{\nu((\hat{o},o))+\epsilon}\mid o\in\mathcal{A}\right\}$triangleright Energy weights
9:         $\mathcal{C}\leftarrow\{o\in\mathcal{A}\mid\nu((\hat{o},o))=0\}$triangleright New coverage
@@
11:              $o^{*}\leftarrow arg\,max_{o\in\mathcal{C}}\phi(o,S)$triangleright $\phi$ indicates pair transition coverage
@@
13:              $o^{*}\sim P(o)\propto\mathcal{E}$triangleright Energy-based sampling
@@
15:         $\nu\leftarrow\nu((\hat{o},o^{*}))+1$triangleright Update frequency recorder
16:         $S\leftarrow\delta(S,o^{*})$triangleright Update state schema
17:         $O\leftarrow O\circ o^{*}$triangleright Append to trace
@@
By explicitly modelling both the state schema and the transitions, TraceGenerator ensures that the generated API sequences are valid and executable. The generation process is outlined in Algorithm 1. Initially, TraceGenerator starts by creating a collection of random variables to serve as the initial states, which could include user-defined parameters or data items from a remote server (line 2). As the sequence generation progresses, at every step $i$, TraceGenerator determines the subsequent API call by evaluating all potential transitions. For each transition, it checks the compatibility of the current variable space with the possible API calls, ensuring that only feasible API calls are considered.
To further improve the quality and diversity of the generated sequences, we employ energy-based sampling [48] strategies when selecting a transition from the candidate pool. This approach is guided by the pair transition; that is, by treating the current transition as a “consumer,” we collect all states associated with its input parameters (e.g., “producers”) and identify the most recent transitions that affected these states. This allows us to construct transition pairs in the format (previous transition, current transition) using states as bridges. Our hypothesis is that increasing the diversity of such transition pairs will enable richer and more complex inter-dependency relationships in generated traces, thereby initiating a more comprehensive evaluation of LLM’s ability on API usage.

### v2

@@ Abstract
By integrating tools from external APIs, Large Language Models (LLMs) have expanded their promising capabilities in a diverse spectrum of complex real-world tasks. However, testing, evaluation, and analysis of LLM tool use remain in their early stages. Most existing benchmarks rely on manually collected test cases, many of which cannot be automatically checked for rigorous semantic correctness and instead depend on static methods such as string matching. Additionally, these benchmarks often overlook the complex interactions that occur between sequential API calls, which are common in real-world applications. To fill the gap, we introduce StateGen, an automated framework designed to generate diverse coding tasks involving sequential API interactions. StateGen combines state-machine-based API constraint solving and validation, energy-based sampling, and control-flow injection to generate executable programs. These programs are then translated into human-like natural language task descriptions through a collaboration of two LLM agents. Utilizing StateGen, we construct StateEval, a benchmark encompassing 120 manually verified test cases spanning across three representative scenarios: Session Service, Tensor Operation, and ElevenLabs MCP. Experimental results confirm that StateGen can effectively generate challenging and realistic API-oriented tasks, highlighting areas for improvement in current LLMs incorporating APIs. We make our framework and benchmark publicly available at StateEval to support future research.
@@ I Introduction
I-A Motivation
The meteoric advancements in Large Language Models (LLMs) have enabled significant performance leaps across domains [chang2024survey, wang2024survey, naveed2023comprehensive] and have transformed many areas of software engineering [jiang2024survey, liu2023your, ahmed2022few, ahmed2024automatic, yang2024exploring, pan2024lost]. Building on this progress, stakeholders have found that integrating tool usage with LLM-driven systems can further extend their capabilities [patil2024gorilla, le2024kat, qin2023toolllm, li2023api], enabling tasks such as numerical computation [xu2025llm] and access to up-to-date information [rag_survey] that are otherwise difficult for LLMs to perform on their own.
Despite the promising progress of this tool integration paradigm, using LLMs to perform designated tasks with external APIs remains a practical yet challenging problem. In real-world software development, practitioners typically start by analyzing and decomposing the task requirements and then understanding the functionalities and usage of available APIs from their documentation [zhong2017empirical]. Eventually, multiple APIs need to be managed in an appropriate sequence, with correctly determined inputs and parameters, to achieve the intended functionalities [nguyen2014mining, liu2021learning]. Such a multi-step development pipeline demands advanced abilities in reasoning, management, planning, and tool calling, posing significant challenges for LLMs, which act as the central control unit to coordinate these processes. As illustrated in Figure 1, converting a given text into speech with a specific style demands the LLM to extract key user requirements (i.e., “convert to speech”, “professional feminine voice”) from the input prompt and inspect the available APIs to determine the appropriate ones for execution. Given the numerous API candidates, the LLM must identify the correct one capable of accomplishing the task and manage the parameters and variables accordingly. Considering the complexity involved, it is crucial to understand to what extent LLMs can comprehend the functionalities of various APIs and facilitate sequential API calls to fulfill designated objectives. Therefore, systematic testing and evaluation of such intricate yet realistic scenarios are one of the foundational components for moving the field forward.
@@ I-A Motivation
Several recent benchmarks [liu2023your, chen2021codex, zhuo2025bigcodebench, lai2023ds] have been introduced to systematically evaluate the coding capabilities of LLMs across a variety of contexts. However, most of these efforts primarily focus on general-purpose, small-scale coding tasks, often overlooking scenarios that require reasoning over interdependent API interactions. While some studies [zan2023private, liu2023codegen4libs, wu2024comprehensive] have begun to incorporate API-centric tasks from specific libraries, their evaluation scope is typically confined to a narrow set of libraries, thereby limiting the generalizability of their findings to broader contexts.
@@
On the other hand, existing benchmarks tailored for function calling tasks tend to rely on manually curated datasets, which are labor-intensive to construct and difficult to scale to new domains or task types. Moreover, many related works concentrate on addressing daily user requests that involve only a handful of APIs, for example, rather than tackling more complex programming scenarios that demand sophisticated orchestration of multiple APIs. This gap highlights the need for more systematic and scalable benchmarks that can better reflect the challenges encountered in real-world, API-rich programming environments.
Consequently, a key research question then arises:
@@
How can we design an automated test generation approach that systematically evaluates LLMs’ ability in understanding sequential API calls and managing the associated program states?
I-B Challenges
Addressing this question involves designing a test case [^1] In this study, we use the term test case to refer to a user prompt designed to evaluate an LLM’s ability to generate programs involving sequential API calls. We use the term oracle to denote the mechanism for assessing the correctness of the generated code. generation framework that can overcome several key challenges:
1. Generating Natural Language Inputs. Users typically instruct LLMs with informal, natural language. Consequently, the test engine has to be able to automatically generate such prompts to serve as realistic test inputs.
2. Managing Complex Program States. The code corresponding to the realistic user requests involving sequential API calls often has interdependent relationships (e.g., producer-consumer relationship) across different calls. This requires a stateful test engine that can manage program states across these interactions. However, this introduces a tension: the engine must reconcile the informality and ambiguity of user requests with the strict, state-dependent logic of the program.
3. Automating Correctness Evaluation. While existing methods often rely on an LLM-as-judge or simple string matching [huang-etal-2024-planning-creation], these approaches can be unreliable for evaluating the correctness of generated programs. We argue that a deterministic, test-oracle-based evaluation is essential for a rigorous and accurate assessment.
I-C Contribution
To address the aforementioned challenges, we propose an automated test generation framework StateGen and a benchmark StateEval in this study. The StateGen workflow begins by generating executable programs from existing API documentation. The generation stage, based on a stateful engine, uses API compatibility checking, energy-based API sampling, and control-flow injection to create programs containing sequential API calls, addressing the state management demands in Challenge 2. Next, to fulfill the requirements of Challenge 1, two LLM agents collaborate to translate these programs into coherent natural language instructions. These instructions then serve as the test inputs for the target LLMs. Finally, oracle construction (Challenge 3 ) is performed through a replay mechanism: the generated code is executed, and both its final outputs and side effects are recorded as the test oracle.
Building upon StateGen, we introduce a new benchmark, StateEval, comprising 120 API-oriented test cases across three representative application scenarios. In summary, this paper makes the following contributions:
@@ I-C Contribution
  • Manually Verified Benchmark. Using this framework, we implement three scenarios: RESTful API calling (Session Service [arcuri2019restful]), tensor manipulation (Tensor Operation [pytorch]), and text-to-speech MCP-based LLM tool calling (ElevenLabs MCP [elevenlabs_mcp]). We generate and manually verify 120 test cases, forming a new benchmark, StateEval, to support future evaluation and advancement of LLM-driven API usage.
@@
  • Artifacts. We publicly release our artifacts, including the implementation of StateGen (approximately 7,800 lines of code) and our benchmark, StateEval (120 manually verified test cases), via a GitHub repository: StateEval.
II Background
In this study, we define LLM tool usage as the generation of code involving function calls to accomplish a task. This definition spans two scenarios. The first (Section II-A) lies in the software engineering context, where LLMs are instructed to produce code with API calls based on engineers’ requirements (e.g., create a new object using a POST request). The second (Section II-B) concerns enhancing LLM capabilities by integrating diverse external tools as callable functions, enabling LLMs to address general user requests more effectively (e.g., book a flight this Wednesday). The former is more classical and engineering-oriented (e.g., CRUD operations), whereas the latter reflects emerging paradigms and standards (e.g., MCP [anthropic_mcp]). Despite these differences, we treat both as instances of LLM tool usage and model them within a unified framework.
II-A LLM in Software Engineering
Modern software development is increasingly facilitated by LLMs that can (1) synthesize code from natural language and (2) leverage external tools and APIs to accomplish complex tasks. These two capabilities are tightly coupled in practice. That is, code generation determines how an LLM implements an algorithm or workflow, and tool usage dictates to what extent an LLM can utilize the external functionality and comprehend the hidden sequential workflows. Effective evaluation, therefore, needs to capture not only function-level correctness but also stateful, sequential behaviors, as well as data and control dependencies when multiple APIs are involved.
Existing test paradigms for LLM code generation provide useful building blocks [chen2021codex, liu2023your, austin2021program, jimenez2023swe, zhuo2025bigcodebench] (e.g., unit tests, executable oracles, and string matching), but they often focus on general coding or engineering problems instead of sequential API calls with multi-step interdependencies.
II-B LLM Enhanced with Tool
By empowering LLMs to use external tools through APIs, we can extend their capabilities beyond static, pre-trained knowledge to perform complex, real-world tasks. This paradigm, explored in comprehensive surveys like ToolLLM [qin2023toolllm], shifts the role of an LLM from a text generator to an intelligent agent. Foundational work such as ReAct [yao2023react] demonstrated that LLMs can effectively combine reasoning traces with task-specific actions (i.e., API calls), while Toolformer [schick2023toolformer] showed that LLMs can learn to use tools in a self-supervised manner. This enables them to reason, plan, and act upon their environment to solve problems that require up-to-date information or specialized computations.
A critical challenge in LLM tool usage is the necessity for sequential and stateful API interaction. Most meaningful tasks require an LLM to generate not just one, but a sequence of API calls in a logical order [qin2023toolllm]. For example, as illustrated in our motivating example (Figure 1), converting text to speech with a specific voice may first require a Search() API to find a voice ID, which is then passed to a Txt2Sph() API. This process is inherently stateful: the output of one API call becomes a critical piece of the program’s state that must be correctly managed and used as an input for a subsequent call. This requirement elevates the task from simple tool selection to a form of program synthesis, where the LLM has to manage data dependencies across multiple steps.
To mitigate errors from incorrect state handling, structured frameworks are emerging. For instance, StateFlow [wu2024stateflow] models workflows as state machines, but its application has primarily focused on domains like shell (bash) and SQL command execution. This leaves a critical gap for general-purpose API calls, where incorrect state management amplifies the risk of failure, a concern highlighted by safety evaluations like ToolEmu [ruan2023identifying]. The limitations of existing research and the high stakes involved underscore the urgent need for scalable methods to automatically generate diverse and verifiable test cases from API documentation to systematically evaluate stateful tool use.
@@ III Approach
III-A Task Definition of LLM Under Test
@@
We begin by formally defining the problem. An LLM under test is asked to solve a state-dependent code generation task. More specifically, let $S=(s_{1},s_{2},\ldots,s_{n})$ denote the set of all program states, where each $s_{j}$ corresponds to its respective $j$-th state. A given state may represent either observable program status (e.g., locally defined variables) or external elements (e.g., the data items stored in remote databases). Given the natural language instructions specifying the desired tasks for a sequence of API calls, along with documentation that defines the usages of corresponding APIs, the objective for LLMs is to generate code that is syntactically and semantically correct in performing the intended functionalities. The generated code should encompass a sequence of $n$ API calls $O=(o_{1},o_{2},\ldots,o_{n})$, where $o_{j}$ denotes the API call at the $j$-th timestamp and can have side effects on states in $S$.
@@
To evaluate the capabilities of LLMs in addressing such intricate API-oriented coding tasks, we propose StateGen, an automated test case generation framework for LLMs with sequential API calls. The overall workflow is shown in Fig 2. At a high level, StateGen follows a reverse-generation strategy, beginning with generating valid, executable sequences of API calls of a given length (Section III-C). These traces are then used as a foundation to construct more intricate executable programs by incorporating control flows, thereby enhancing diversity and complexity (Section III-D). Next, StateGen employs an LLM-driven multi-agent translation process (Section III-E) to transform these programs into human-like natural language instructions. To obtain test oracles for automated evaluation, StateGen executes the generated programs in a local environment and records the state transitions throughout program execution (Section III-F). These captured transitions serve as ground truth for subsequent evaluation.
III-B Scenario Selection
Before presenting our methodology, we highlight a central design decision that shapes our work: instead of collecting existing API traces from real-world systems, as most prior studies do, we build a stateful fuzzing engine that explicitly models how each API behaves within a scenario. Existing evaluations of LLM tool use tend to emphasize breadth by constructing large API datasets through crawling or repository mining [basu2025nestful, wu2024comprehensive, saha2024sequential]. Although these datasets cover many tools, they usually treat API calls as isolated or short sequences and often rely on static text matching without execution guarantees. As a result, they overlook the combinatorial complexity of real software, where the challenge lies not only in choosing the right API but also in orchestrating long chains of dependent operations.
To address this limitation, we depart from the “massive-scale” strategy and instead adopt a depth-first fuzzing approach. By formally modeling a focused set of domains with state machines, we open a large space of valid execution paths (exponentially growing with respect to sequence length). This enables us to generate long, state-aware API sequences that are structurally sound yet semantically challenging. This dynamic generation paradigm also reduces the risk of data contamination, which is more difficult to avoid when using static benchmarks that an LLM may have already been exposed to during training.
This design philosophy requires us to carefully select which scenarios to model and fuzz. To reflect the range of challenges seen in real-world function calling, we select three usage scenarios that span distinct stages of technological adoption and specialization. Our selection ranges from mature, general-purpose web standards (such as RESTful APIs) to specialized scientific libraries and emerging LLM-agent-based protocols. This diversity enables us to assess the abilities of LLMs under test across technologies with widely varying levels of abstraction and constraint rigidity, ranging from the explicit specifications of classical web services to the implicit mathematical requirements of deep learning and the schema-driven interfaces of future tool-use systems.
  • Session Service: This application represents the foundational layer of software infrastructure. We utilize the Session Service from the EvoMaster Benchmark (EMB) [arcuri2019restful]. As a proxy for a traditional web application, this service manages user sessions via MongoDB using standard RESTful patterns and OpenAPI specifications [openapi]. By targeting this ubiquitous domain, we validate the LLM’s ability to handle stateful resource management within the stable, well-established standards that characterize the majority of existing web APIs. For this scenario, we selected five APIs.
  • PyTorch Tensor Operation: This application represents the domain-specific high-performance computing sector. We select PyTorch [pytorch], a widely used industry-level deep learning library. Different from general-purpose web services, this domain requires compliance with underlying constraints related to data dimensionality and type handling. LLMs’s abilities to track strict mathematical constraints and maintain API compatibility in high-dimensional operations will be evaluated in this application. We select six APIs for this scenario: four related to tensor manipulation and two focused on tensor computation (Linear and Conv2d). Related documentation was sourced directly from the official website.
  • ElevenLabs MCP: Model Context Protocol (MCP) [anthropic_mcp] represents the nascent frontier of automated LLLM agent-to-tool interaction. We adopt an implementation from ElevenLabs [elevenlabs_mcp]. This implementation offers a set of AI tools for speech processing, such as text-to-speech, transcription, and outbound phone calls. As a newly introduced standard designed explicitly for LLM-tool interoperability, MCP differs significantly from legacy APIs in its intention and structure. This scenario evaluates the framework’s adaptability to standardized agent interfaces, ensuring that our generation pipeline remains robust when applied to cutting-edge protocols with limited historical training data. From this collection, we choose six APIs.
These three scenarios span conventional RESTful services, data‑intensive scientific computing, and recent LLM‑driven tool use, covering 17 APIs. Together, they provide diverse, realistic settings to evaluate how LLMs handle API calls and to assess StateGen. Our choice to orchestrate 5 to 6 distinct APIs per task is also supported by a large empirical study [zhong2017empirical] of millions of lines of code from popular open‑source projects (e.g., ZooKeeper), which finds that most methods contain short API sequences. Coordinating 5 to 6 APIs can already capture the complexity of most real-world single‑method workflows. Furthermore, StateGen remains extensible: additional APIs can be included whenever they can be modeled with our state machine.
For clarity, when illustrating our methodology, we will use a simplified PyTorch Tensor Operation example to illustrate the idea.
III-C Trace Generation
@@
Figure 2: Workflow overview of StateGen. Trace Generation (Sec III-C) forms the backbone of StateGen, producing valid API sequences through state machines while ensuring diversity with energy-based sampling. Program Generation (Sec III-D) then assembles these traces with appropriate initialization and control flow structures. Finally, Instruction Translation (Sec III-E) employs a multi-agent system to convert the generated programs back into natural language descriptions for evaluation.
@@
Generating valid sequential API calls poses significant technical challenges in this work, as it requires precise modelling of program states and meticulous selection of valid API calls at each step. Traditional approaches that address this problem have been extensively studied in the context of web service testing, e.g., RESTful API testing [arcuri2019restful, atlidakis2019restler, viglianisi2020resttestgen, martin2021restest, kim2022automated, liu2022morest, golmohammadi2023testing]. However, these methods are difficult to apply directly, as they focus primarily on CRUD operations, which represent only a subset of our applications. Adapting them for general LLM tool calling would require significant additional effort.
@@
Despite this, the core concept remains unchanged. Prior work in RESTful API testing can generally be categorized into top-down [viglianisi2020resttestgen] and bottom-up [atlidakis2019restler, liu2022morest] strategies. In top-down approaches, the overall API call graph is built in advance, and each API call is instantiated according to the graph. Conversely, bottom-up approaches construct the API sequence progressively by adding one call at a time. In this work, we adopt a bottom-up strategy, allowing us to employ a “lazy evaluation” strategy for determining API inputs and parameters. Specifically, we compute only the valid and relevant search space at each step of API selection rather than pre-solving constraints for the entire API call graph, which otherwise would demand substantial computational efforts on optimization to attain practical efficiency [liu2023nnsmith].
@@
$o:\mathbb{S}\times\mathbb{P}\rightarrow\mathbb{S}],\quad\forall o\in\mathbb{O}$ (1)
@@
[…]
2:$O$ $\triangleright$ Generated trace of API calls
@@
4:  $O\leftarrow\emptyset$, $S\leftarrow RandomInit$ $\triangleright$ Initialization
5:  $\hat{o}\leftarrow\bot \triangleright$ Initialize previous transition to None
@@
7:   $\mathcal{A}\leftarrow\{o\in\mathcal{O}\mid\text{valid}(o,S,O)\}$ $\triangleright$ Obtain transitions
8:   $\mathcal{E}\leftarrow\left\{\frac{1}{\nu((\hat{o},o))+\epsilon}\mid o\in\mathcal{A}\right\}$ $\triangleright$ Energy weights
9:   $\mathcal{C}\leftarrow\{o\in\mathcal{A}\mid\nu((\hat{o},o))=0\}$ $\triangleright$ New coverage
@@
11:     $o^{*}\leftarrow arg\,max_{o\in\mathcal{C}}\phi(o,S)$ $\triangleright$ $\phi$ indicates pair transition coverage
@@
13:     $o^{*}\sim P(o)\propto\mathcal{E}$ $\triangleright$ Energy-based sampling
@@
15:   $\nu\leftarrow\nu((\hat{o},o^{*}))+1$ $\triangleright$ Update frequency recorder
16:   $S\leftarrow\delta(S,o^{*})$ $\triangleright$ Update state schema
17:   $O\leftarrow O\circ o^{*}$ $\triangleright$ Append to trace
@@
For illustration, we take the PyTorch Tensor Operation as an example. Consider a step in the trace generation process where we have generated the first API call and are about to generate the second (Code Snippet 1).
Code Snippet 1: First API
    1  ut1 = torch.load("user1")  2  ut2 = torch.load("user2")  3 # API 1: concatenate the tensors from 4 # initilized variables ut1 and ut2  5 # along dimension 0 6  r_1 = torch.cat((ut1, ut2), 0)
At this step, the TraceGenerator records the state of all intermediate variables, including the shapes and values of tensors ut1, ut2, and r_1. Furthermore, we update auxiliary information associated with each variable, such as recording that both ut1 and ut2 now participate in a torch.cat operation. In addition, r_1 is marked as a newly created variable produced by this function call and has not yet been used as an input to any subsequent operation. Recording such operation history helps the fuzzer track interdependencies between variables, which is essential for selecting the next API call.
By explicitly modelling both the state schema and the transitions, TraceGenerator ensures that the generated API sequences are valid and executable. The generation process is outlined in Algorithm 1. Initially, TraceGenerator starts by creating a collection of random variables to serve as the initial states, which could include user-defined parameters or data items from a remote server (line 2). Note that once the fuzzer finalizes the values of these initialized variables, they are stored persistently on disk and later loaded during evaluation, just as shown in the first two lines of the Snippet 1. This is essential to avoid randomness in the evaluation.
As the sequence generation progresses, at every step $i$, TraceGenerator determines the subsequent API call by iterating through all available APIs and checks whether each one is a feasible next candidate. For example, the conv2d operation in our Pytorch Tensor Operation example applies a 2D convolution to an input tensor using a weight tensor, with parameters such as stride, padding, and dilation. The TraceGenerator checks whether any valid parameter configuration exists that makes this operation legal, given the shapes and types of the existing tensors that can serve as input and weight. This constraint checking and parameter solving occur on the fly at every step of generation. For Pytorch Tensor Operation, parameter selection is guided by the mathematical rules behind each operation—that is, solving the corresponding equations to determine feasible parameter values. In other scenarios, this process involves checking whether the remote database contains the necessary data items to support a function call, or verifying that the objects currently in scope can serve as valid inputs for the relevant MCP interface.
To further improve the quality and diversity of the generated sequences, we employ energy-based sampling [bohme2016coverage] strategies when selecting a transition from the candidate pool. This approach is guided by the pair transition; that is, by treating the current transition as a “consumer,” we collect all states associated with its input parameters (e.g., “producers”) and identify the most recent transitions that affected these states. This allows us to construct transition pairs in the format (previous transition, current transition) using states as bridges. Our hypothesis is that increasing the diversity of such transition pairs will enable richer and more complex inter-dependency relationships in generated traces, thereby initiating a more comprehensive evaluation of LLM’s ability on API usage.